Solve the (4x3) game with pay off matrix.


8 5 8

[A] = 8 6 5

7 4 5

6 5 6


 At each stage, clearly explain the steps involved.


1
Expert's answer
2021-06-07T16:35:40-0400

If all the elements of Column-i are greater than or equal to the corresponding elements of any other Column-j, then the Column-i is dominated by the Column-j and it is removed from the matrix

So, we can remove 1st column (col1 "\ge" col2)

 If all the elements of a Row-i are less than or equal to the corresponding elements of any other Row-j, then the Row-i is dominated by the Row-j and it is removed from the matrix.

So, we can remove 3rd row (row3 "\le" row2), and 4th row (row4 "\le" row1).


Then we get matrix:

5 8

8 6


This game has no saddle point.


Then:

A play’s (p1,p2):

"p_1=\frac{d-c}{(a+d)-(b+c)}" , "p_2=1-p_1"


B play’s (q1,q2):

"q_1=\frac{d-b}{(a+d)-(b+c)}" , "q_2=1-q_1"


Value of the game:

"V=\frac{ad-bc}{(a+d)-(b+c)}"


From payoff matrix

"\begin{pmatrix}
 a & b \\
 c & d
\end{pmatrix}"

we have:

"a=5,b=8,c=8,d=6"


"p_1=\frac{6-8}{(5+6)-(8+8)}=\frac{2}{5}" , p2=3/5


"q_1=\frac{6-8}{(5+6)-(8+8)}=\frac{2}{5}" , q2=3/5


"V=\frac{5\cdot6-8\cdot8}{(5+6)-(8+8)}=\frac{34}{5}"